# Compositional amortized inference for large-scale hierarchical Bayesian models

### v1

@@ Acknowledgments
J.A.: Conceptualization, Methodology, Software, Formal analysis, Validation, Visualization, Funding acquisition, Writing – original draft, Writing – review & editing. V.P.: Data curation, Software, Writing – review & editing. C.S.: Investigation, Data curation, Writing – review & editing. M.B.: Supervision, Investigation, Data curation, Writing – review & editing. X.I.: Supervision, Project administration, Resources, Writing – review & editing. J.H.: Supervision, Project administration, Funding acquisition, Writing – review & editing. S.T.R.: Conceptualization, Methodology, Supervision, Resources, Funding acquisition, Writing – original draft, Writing – review & editing.
@@ A.2 Mini-batch estimator is unbiased
$\hat{s}_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y},\lambda_{t})=(1-J)(1-t)\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{\theta})+\frac{J}{M}\sum_{j=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t})$
@@
$\mathbb{E}\left[\frac{J}{M}\sum_{j=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t})\right]=\frac{J}{M}\sum_{j=1}^{M}\mathbb{E}_{\mathbf{Y}_{j}}\left[s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t})\right].$
@@
$\mathbb{E}_{\mathbf{Y}_{j}}\left[s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t})\right]=\frac{1}{J}\sum_{j=1}^{J}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t}),$
@@
$\mathbb{E}\left[\frac{J}{M}\sum_{j=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t})\right]=\frac{J}{M}\cdot M\cdot\frac{1}{J}\sum_{j=1}^{J}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t})=\sum_{j=1}^{J}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j},\lambda_{t}).$
Adding the constant prior term $(1-J)(1-t)\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{\theta})$ yields the full compositional score. Hence, the estimator is unbiased. ∎
@@ A.3 Evaluation metrics
RMSE measures the deviation between posterior samples and the ground-truth parameters. Given posterior samples $\hat{\boldsymbol{\theta}}^{(s)}_{ij}$ for parameters $j$ in dataset $i$, and true parameters $\boldsymbol{\theta}_{ij}$, the RMSE is defined as:
@@
$\mathrm{KL}(q\,\|\,p)=\frac{1}{2}\left[\log\frac{|\mathbf{\Sigma}_{p}|}{|\mathbf{\Sigma}_{q}|}-d+\mathrm{Tr}(\mathbf{\Sigma}_{p}^{-1}\mathbf{\Sigma}_{q})+(\mathbf{\mu}_{q}-\mathbf{\mu}_{p})^{\top}\mathbf{\Sigma}_{p}^{-1}(\mathbf{\mu}_{q}-\mathbf{\mu}_{p})\right],$
where $\mathbf{\mu}_{q}$, $\mathbf{\Sigma}_{q}$ are the empirical mean and covariance of posterior samples, and $\mathbf{\mu}_{p}$, $\mathbf{\Sigma}_{p}$ are the parameters of the analytical posterior.
@@ A.4 Score model architectures & training
    • Final representations from both paths are concatenated to produce a summary embedding which are then projected by a linear layer to a fixed summary dimension of size 18.
We parameterize our score models to predict the more stable $\hat{\mathbf{v}}_{t}:=\alpha_{t}\boldsymbol{\epsilon}-\sigma_{t}\boldsymbol{\theta}_{t}$, and then transform the output to $\hat{\boldsymbol{\epsilon}}_{t}$, as it has been shown that this parameterization is more stable for all $t$, whereas noise-prediction becomes harder for $t$ close to 0 where the signal increases and noise decreases Salimans and Ho [2022]. Furthermore, we condition the score network on the signal-to-noise ratio (SNR), normalized to the interval $[-1,1]$ similar to the preconditioning introduced in [Karras et al., 2022]. The data and parameters are always standardized, and the prior scores are adjusted accordingly by multiplying them by the standard deviation of the parameters.
@@
  • Cosine schedule by Nichol and Dhariwal [2021] (with $s{=}0$ during training)
@@
All our noise schedules are truncated such that the log signal-to-noise ratio is $\lambda_{t}\in[-15,15]$ to avoid instabilities in sampling as detailed in [Kingma and Gao, 2023]. For the EDM schedule we set $\lambda_{t}\in[-\log\sigma^{2}_{\text{max}},-\log\sigma^{2}_{\text{min}}]$ as in the original paper.
As weighting function for the loss, we employed the likelihood weighting $w_{t}=g(t)^{2}/\sigma^{2}$ proposed by Song et al. [2021] for the linear and cosine schedules and the original EDM weighting $w_{t}=\exp(-\lambda_{t})+1$ for the EDM schedule [Karras et al., 2022].
@@
We trained all models using AdamW with a cosine annealing learning rate schedule. The initial learning rate is set to $5{\times}10^{-4}$. Models are trained for 1000 epochs on the Gaussian toy example and 3000 epochs on all other settings. In each epoch we generate 10,000 new training samples on the fly. Only for the FLI application we used 30,000 samples per epoch as we found that more training data was needed.
All models were trained on a high-performance computing cluster using a AMD EPYC "Milan" CPU (2.00 GHz), 100 GB DDR4 3200 MHz RAM, and a NVIDIA A40 GPU with 48 GB of memory. Each experiment required 1-2 days for all repeated runs on a high-performance computing infrastructure with up to 50 parallel jobs.
@@
For our experiments, we used the adaptive second-order sampler with maximal 10,000 iterations and the default settings proposed by Jolicoeur-Martineau et al. [2021]. Specifically, we set the absolute error tolerance to $e_{\text{abs}}=0.01\times\text{the number of parameters }$ and the relative tolerance to $e_{\text{rel}}=0.5$. To solve the probability ODE, we used an Euler scheme. For annealed Langevin dynamics, we followed the setup from Geffner et al. [2023], using 5 Langevin steps per iteration, a maximum of 2000 iterations, and a step size factor of 0.1.
To find the optimal damping factor $d_{1}$ and optimal shift $s$ for a certain task, we run a hyper-parameter search optimizing the combined loss of RMSE and calibration error on a test data set. We also considered $d_{0}<1$, and found that sometimes this can improve accuracy of the posterior, even though the true posterior is theoretically not recovered.
Code
The software code and data for the experiments is available in a public GitHub repository: https://anonymous.4open.science/r/hierarchical-abi-submission-7409/.
@@ A.5 Experiment 1: Gaussian toy model
Model definition:
$\boldsymbol{\theta}\sim\mathcal{N}(\boldsymbol{\theta},\sigma\mathbf{I})$
with $\sigma=0.1$ and dimension set to 10. Given a normal prior, with the same scale, the posterior is also Gaussian and we can calculate it analytically. Here, we do not employ a summary network.
@@
Figure 7: Evaluation of the error-damping estimator for the toy model. Different evaluation metrics are shown for different mini-batch sizes or varying numbers of subsets of groups. For each experiment 10 runs were performed. The median and median absolute deviation is reported besides for those runs where none converged.
@@
Figure 8: Evaluation of the different noise schedules for the toy model. For each experiment 10 runs were performed. The median and median absolute deviation is reported besides for those runs where none converged. Both methods fail for already 1,000 groups, where the standard cosine schedule still converges.
@@ A.6 Experiment 2: hierarchical AR(1) model
$\tilde{\boldsymbol{\eta}}_{j}\sim\mathcal{N}(0,\sigma\mathbf{I}),\qquad\boldsymbol{\eta}_{j}=2\operatorname{sigmoid}(\beta+\tilde{\boldsymbol{\eta}}_{j})-1.$
@@
$\mathbf{Y}_{j,0}$ $\sim\mathcal{N}(0,0.1\mathbf{I})$
$\mathbf{Y}_{j,t}$ $\sim\mathcal{N}(\alpha+\boldsymbol{\eta}_{j}\mathbf{Y}_{j,t-1},0.1\mathbf{I}),\quad t=1,\ldots T-1.$
On the local level, we perform inference on $\tilde{\boldsymbol{\eta}}$ and afterwards transform $\tilde{\boldsymbol{\eta}}$ to $\boldsymbol{\eta}$ as NUTS [as implemented in Stan Carpenter et al., 2017] performs better on non-centered parameterizations [Betancourt and Girolami, 2015]. Here, we do not employ a summary network.
(a) Varying mini-batch sizes. Convergence is achieved only for the smallest data set.
(b) Varying cosine shifts $s$.
(c) Varying number of subsets of groups during training (score trained with a Deep Set).
Figure 9: Evaluation of the error-damping estimator for the hierarchical AR(1) model. For each experiment 10 runs were performed. The median and median absolute deviation is reported besides for those runs where none converged. A mini-batch size of 10% of the data is employed and score models are trained on a single group.
(a) NUTS: Recovery of global parameters.
(b) NUTS: Simulation-based calibration diagnostics.
Figure 10: Calibration verification of NUTS. NUTS, regarded as a gold standard for Bayesian inference, can recover global parameters and is well calibrated for the $32{\times}32$ grid. However, for larger grids the computational costs explode.
@@ A.7 Experiment 3: fluorescence lifetime imaging (FLI) model
The traditional ways of fitting this kind of models are reviewed in Torrado et al. [2024].
@@
Alternatively, we could have used a non-centered parameterization commonly found in the mixed-effects literature (such as for the hierarchical AR(1) model), where global means are treated as fixed effects and local parameters as additive random effects with a mean of zero and the same standard deviation as described above.
@@
AU565 (HER2+ human breast carcinoma) cells, incubated for 24h with 20 $\mu$g/mL TZM-Alexa Fluor 700 (Donor, D) and 40 $\mu$g/mL TZM-Alexa Fluor 750 (Acceptor, A), were imaged using Förster resonance energy transfer (FRET) microscopy to quantify trastuzumab (TZM) binding. AU565 cells exhibit relative low level of HER2 heterodimerization that correlate with reduced TZM uptake and sensitivity, which is also influenced by culture conditions (2D vs 3D). FLI-FRET analysis allows for the quantification of these dimerization-dependent variations in live cells by assessing the proximity of donor and acceptor-labeled TZM.
@@
Figure 12: Assessing inference of local parameters for the FLI model. Synthetic data on a $32{\times}32$ grid was generated. We compare our hierarchical approach against the standard non-hierarchical pixel-wise MLE.
@@
(c) Correlation of $\tau_{\text{mean}}$ estimates. Our local posterior medians vs. pixel-wise MLE estimates.
Figure 13: Assessing inference of local parameters for the FLI model on real data. We compare our hierarchical approach against the standard non-hierarchical pixel-wise MLE. Due to low photon count, the average lifetime $\tau^{\text{mean}}$ is the most reliable quantity for this non-hierarchical method.

### v3

@@ Acknowledgments
J.A.: Conceptualization, Methodology, Software, Formal analysis, Validation, Visualization, Funding acquisition, Writing – original draft, Writing – review & editing. V.P.: Data curation, Software, Writing – review & editing. C.S.: Investigation, Data curation, Writing – review & editing. M.B.: Supervision, Investigation, Data curation, Writing – review & editing. X.I.: Supervision, Project administration, Resources, Writing – review & editing. J.H.: Supervision, Project administration, Funding acquisition, Writing – review & editing. S.T.R.: Conceptualization, Methodology, Software, Supervision, Resources, Funding acquisition, Writing – original draft, Writing – review & editing.
@@ A.2 Mini-batch estimator is unbiased
$\hat{s}_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y},\lambda_{t})=(1-J)(1-t)\nabla_{\eta_{t}b}\log p(\boldsymbol{\eta}_{t})+\frac{J}{M}\sum_{j=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t})$
@@
$\mathbb{E}\left[\frac{J}{M}\sum_{j=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t})\right]=\frac{J}{M}\sum_{j=1}^{M}\mathbb{E}_{\mathbf{Y}_{j}}\left[s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t})\right].$
@@
$\mathbb{E}_{\mathbf{Y}_{j}}\left[s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t})\right]=\frac{1}{J}\sum_{j=1}^{J}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t}),$
@@
$\mathbb{E}\left[\frac{J}{M}\sum_{j=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t})\right]=\frac{J}{M}\cdot M\cdot\frac{1}{J}\sum_{j=1}^{J}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t})=\sum_{j=1}^{J}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j},\lambda_{t}).$
Adding the constant prior term $(1-J)(1-t)\nabla_{\boldsymbol{\eta}}\log p(\boldsymbol{\eta}_{t})$ yields the full compositional score. Hence, the estimator is unbiased. ∎
@@ A.3 Evaluation metrics
RMSE measures the deviation between posterior samples and the ground-truth parameters. Given posterior samples $\hat{\boldsymbol{\theta}}^{(s)}_{ij}$ (local or global) for parameters $j$ in the dataset $i$, and true parameters $\boldsymbol{\theta}_{ij}$, the RMSE is defined as:
@@
$\mathrm{KL}(q\,\|\,p)=\frac{1}{2}\left[\log\frac{|\boldsymbol{\Sigma}_{p}|}{|\boldsymbol{\Sigma}_{q}|}-d+\mathrm{Tr}(\boldsymbol{\Sigma}_{p}^{-1}\boldsymbol{\Sigma}_{q})+(\boldsymbol{\mu}_{q}-\boldsymbol{\mu}_{p})^{\top}\boldsymbol{\Sigma}_{p}^{-1}(\boldsymbol{\mu}_{q}-\boldsymbol{\mu}_{p})\right],$
where $\boldsymbol{\mu}_{q}$, $\boldsymbol{\Sigma}_{q}$ are the empirical mean and covariance of posterior samples, and $\boldsymbol{\mu}_{p}$, $\boldsymbol{\Sigma}_{p}$ are the parameters of the analytical posterior.
Posterior predictive $p$-value.
The posterior predictive $p$-value evaluates how well the observed data are covered by the posterior predictive distribution. In a well‑specified model, these pp‑values are approximately uniform on $[0,1]$; thus, their expectation should be approximately 0.5. For $S$ posterior samples, let
$f_{t}(\boldsymbol{\theta})$ $=\operatorname{median}\left(\{y_{t}^{\text{rep},(s)}\sim p(\mathbf{Y}\mid\boldsymbol{\theta})\}^{S}_{s=1}\right),\quad\widehat{\operatorname{Var}}_{t}(\boldsymbol{\theta})=\frac{1}{S-1}\sum_{s=1}^{S}\left(y_{t}^{\text{rep},(s)}-f_{t}(\boldsymbol{\theta})\right)^{2}.$
For each posterior draw $\boldsymbol{\theta}^{(s)}$, define the discrepancy as
$D\left(\mathbf{y},\boldsymbol{\theta}\right)$ $=\sum_{t=1}^{T}\frac{\left(y_{t}-f_{t}(\boldsymbol{\theta})\right)^{2}}{\widehat{\operatorname{Var}}_{t}(\boldsymbol{\theta})},$
and then posterior predictive pp‑value is
$p_{\text{PPC}}=\frac{1}{S}\sum_{s=1}^{S}\mathbf{1}\left(D(\mathbf{y}^{\text{rep},(s)},\boldsymbol{\theta})\geq D(\mathbf{y}_{\text{obs}},\boldsymbol{\theta})\right).$
RMSE, calibration error, posterior contraction and empirical CDFs plots are computed using the diagnostics provided in the BayesFlow toolbox [Radev et al., 2023].
@@ A.4 Score model architectures & training
    • Final representations from both paths are concatenated to produce a summary embedding, which are then projected by a linear layer to a fixed summary dimension of size 18.
We parameterize our score models to predict the more stable $\hat{\mathbf{v}}_{t}:=\alpha_{t}\boldsymbol{\epsilon}-\sigma_{t}\boldsymbol{\theta}_{t}$, and then transform the output to $\hat{\boldsymbol{\epsilon}}_{t}$, as it has been shown that this parameterization is more stable for all $t$, whereas noise-prediction becomes harder for $t$ close to 0 where the signal increases and noise decreases Salimans and Ho [2022]. Furthermore, we condition the score network on the signal-to-noise ratio (SNR), normalized to the interval [1,1][-1,1] similar to the preconditioning introduced in [Karras et al., 2022]. The data and parameters are always standardized, and the prior scores are adjusted accordingly by multiplying them by the standard deviation of the parameters.
@@
  • Cosine schedule by Nichol and Dhariwal [2021] (with s=0s{=}0 during training)
@@
All our noise schedules are truncated such that the log signal-to-noise ratio is $\lambda_{t}\in[-15,15]$ to avoid instabilities in sampling as detailed in [Kingma and Gao, 2023]. For the EDM schedule, we set $\lambda_{t}\in[-\log\sigma^{2}_{\text{max}},-\log\sigma^{2}_{\text{min}}]$ as in the original paper.
As the weighting function for the loss, we employed the likelihood weighting $w_{t}=g(t)^{2}/\sigma^{2}$ proposed by Song et al. [2021] for the linear and cosine schedules and the original EDM weighting $w_{t}=\exp(-\lambda_{t})+1$ for the EDM schedule [Karras et al., 2022].
@@
We trained all models using AdamW with a cosine annealing learning rate schedule. The initial learning rate is set to $5{\times}10^{-4}$. Models are trained for 1000 epochs on the Gaussian toy example and 3000 epochs on all other settings. In each epoch, we generate 10,000 new training samples on the fly, as simulations are cheap. Only for the FLI application we used 30,000 samples per epoch as we found that more training data was needed. However, we found similar performance of our score models when trained with a fixed simulation budget without generating new samples in each epoch. For reference, training a single score estimator for the FLI task completes in 7.6 hours on a single GPU, while for the AR(1) model it takes 0.83 hours.
All models were trained on a high-performance computing cluster using an AMD EPYC "Milan" CPU (2.00 GHz), 100 GB DDR4 3200 MHz RAM, and an NVIDIA A40 GPU with 48 GB of memory. Each experiment required 1–2 days for all repeated runs on a high-performance computing infrastructure with up to 50 parallel jobs.
@@
For our experiments, we used the adaptive second-order sampler with maximal 10,000 iterations and the default settings proposed by Jolicoeur-Martineau et al. [2021]. Specifically, we set the absolute error tolerance to $e_{\text{abs}}=0.01\times\text{the number of parameters }$ and the relative tolerance to $e_{\text{rel}}=0.5$. To solve the probability ODE, we used an Euler scheme. For annealed Langevin dynamics, we followed the setup from Geffner et al. [2023], using 5 Langevin steps per iteration, a maximum of 2000 iterations, and a step size factor of 0.1. For GAUSS, we used the implementation provided by the sbi toolbox [Boelts et al., 2025] with the same diffusion model and training settings as in our own implementation.
To find the optimal damping factor $d_{1}$ and shift $s$ for a certain task, we ran Bayesian optimization with optuna [Akiba et al., 2019] using the sum of the average RMSE and expected calibration error as an optimization criterion. We used search grids $s\in[0,4]$ and $d_{1}\in[1\times 10^{-5},0.1]$. We chose this simple criterion because the hierarchical structure and shrinkage effects in our experiments encourage unimodal behavior by borrowing strength across observations. More expressive criteria can be used in cases where the posteriors exhibit multiple modes. We also considered $d_{0}<1$, and found that this can sometimes improve RMSE and calibration.
@@ A.5 Experiment 1: Gaussian toy model
We define the Gaussian toy model as follows:
$\mathbf{Y}_{i}\sim\mathcal{N}(\boldsymbol{\eta}\mid\sigma^{2}\mathbf{I})$
with $\sigma=0.1$ and $\boldsymbol{\eta}\in\mathbb{R}^{10}$. We observe $\{\mathbf{Y}_{j}\}_{j=1}^{J}$ with varying $J$ and compute the posterior $p(\boldsymbol{\eta}\mid\{\mathbf{Y}_{i}\}_{j=1}^{J})$. Given a normal prior for $\boldsymbol{\eta}$, $\boldsymbol{\eta}\sim\mathcal{N}(\mathbf{0}\mid\sigma^{2}\mathbf{I})$, the posterior is also Gaussian, and we can calculate it analytically:
$p(\boldsymbol{\eta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})\propto\exp\left(-\tfrac{1}{2}\,(\boldsymbol{\eta}-\boldsymbol{\mu}_{J})^{\top}\boldsymbol{\Sigma}_{J}^{-1}(\boldsymbol{\eta}-\boldsymbol{\mu}_{J})\right),$
where $\boldsymbol{\mu}_{J}=\frac{1}{J+1}\sum_{j=1}^{J}\mathbf{Y}_{j}$ and $\boldsymbol{\Sigma}_{J}^{-1}=\frac{J+1}{\sigma^{2}}\mathbf{I}$. Here, we do not employ a summary network.
@@
Figure 6: Evaluation of the error-damping estimator for the toy model. Different evaluation metrics are shown for different mini-batch sizes or varying numbers of subsets of groups. For each experiment, 10 runs were performed. The median and median absolute deviation is reported, besides for those runs where none converged.
@@
Figure 7: Evaluation of the different noise schedules for the toy model. For each experiment, 10 runs were performed. The median and median absolute deviation is reported, besides for those runs where none converged. Both methods fail for already 1,000 groups, where the standard cosine schedule still converges.
@@ A.6 Experiment 2: hierarchical AR(1) model
$\tilde{\boldsymbol{\theta}}_{j}\sim\mathcal{N}(0,\sigma\mathbf{I}),\qquad\boldsymbol{\theta}_{j}=2\operatorname{sigmoid}(\beta+\tilde{\boldsymbol{\theta}}_{j})-1.$
@@
$\mathbf{Y}_{j,0}$ $\sim\mathcal{N}(\mathbf{0},0.1\mathbf{I})$
$\mathbf{Y}_{j,t}$ $\sim\mathcal{N}(\alpha+\boldsymbol{\theta}_{j}\mathbf{Y}_{j,t-1},0.1\mathbf{I}),\quad t=1,\ldots T-1.$
On the local level, we perform inference on $\tilde{\boldsymbol{\theta}}$ and afterward transform $\tilde{\boldsymbol{\theta}}$ to $\boldsymbol{\theta}$ as NUTS [as implemented in Stan Carpenter et al., 2017] performs better on non-centered parameterizations [Betancourt and Girolami, 2015]. Here, we do not employ a summary network.
(a) Varying damping factors $d_{1}$.
(b) Varying mini-batch sizes. Convergence is achieved only for the smallest data set.
(c) Varying cosine shifts $s$.
(d) Varying number of subsets of groups during training (score trained with a Deep Set).
Figure 8: Evaluation of the error-damping estimator for the hierarchical AR(1) model. For each experiment, 10 runs were performed. The median and median absolute deviation is reported, besides for those runs where none converged. A mini-batch size of 10% of the data is employed, and score models are trained on a single group.
@@ A.7 Experiment 3: fluorescence lifetime imaging (FLI) model
The traditional ways of fitting these kinds of models are reviewed in Torrado et al. [2024].
@@
[…]
@@
Figure 10: Assessing inference of local parameters for the FLI model. Synthetic data on a $32{\times}32$ grid was generated. We compared our hierarchical approach against the standard non-hierarchical pixel-wise MLE.
@@
(c) Correlation of $\tau_{\text{mean}}$ estimates.
(d) Local simulations from the hierarchical posterior.
(e) Quality of posterior prediction measured by Bayesian $p$-value.
Figure 11: Assessing inference of local parameters for the FLI model on real data. We compared our hierarchical approach with the standard non-hierarchical pixel-wise MLE. Owing to the low photon count, the average lifetime $\tau^{\text{mean}}$ is the most reliable quantity for this non-hierarchical method. Furthermore, we show additional random simulations from the hierarchical posterior (median and $95\%$ confidence region out of 100 simulations) and a quantitative evaluation of the posterior predictive quality.